OpenStudy (anonymous):

Find the x and y intercepts of the line: 5x + 4y = 20.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well.. as the site says.... set x=0 and then set y= and solve for the other variable

OpenStudy (anonymous):

i still dont know how

OpenStudy (jdoe0001):

well.. you may want to brush up your linear simplification material then

OpenStudy (jdoe0001):

x-intercept 5x + 4(0) = 20. solve for "x" y-intercept 5(0) + 4y = 20. solve for "y"

OpenStudy (anonymous):

wait is it (0,5) (4,0)??

OpenStudy (jdoe0001):

hmm well.. what did you get for "x"?

OpenStudy (anonymous):

those are the only ones I got

OpenStudy (jdoe0001):

well.... yeap , those are it

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

for the x intercept?

OpenStudy (anonymous):

x int its 4,0 y int is 5,0

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

hello? yes??

OpenStudy (jdoe0001):

\(\bf 5x + 4(0) = 20\implies 5x+0=20\implies 5x=20\implies x=\cfrac{\cancel{ 20 }}{\cancel{ 5 }} \\ \quad \\ x=4\qquad {\color{brown}{x-intercept\qquad (4,0)}} \\ \quad \\ 5(0)+4y=20\implies 0+4y=20\implies y=\cfrac{\cancel{ 20 }}{\cancel{ 4 }} \\ \quad \\ y=5\qquad {\color{brown}{ y-intercept\qquad (0,5)}}\)

OpenStudy (anonymous):

ohh ok almost right